A helium-neon laser (?=633nm) illuminates a diffraction grating. The distance between the two m=1bright fringes is 38 cm on a screen 1.6 m behind the grating.

a. What is the spacing between slits of the grating?

Express your answer to two significant figures and include the appropriate units.

d = ???

NOTE: 4.3*10-6m is wrong.
Answer #1

here,

lamda = 633 * 10^-9 m

screen distance , D = 1.6 m

let the slit spacing be d

y = m * lamda * D/d

0.38 = 1 * 633 * 10^-9 * 1.6 /d

solving for d

d = 2.67 * 10^-6 m

the slit spacing is 2.67 * 10^-6 m